OpenStudy (anonymous):

evaluate 7x^3y^-2 for x=1 and y=-2

OpenStudy (mathstudent55):

Substitute x with 1 and y with -2 and do the operations. Just remember the correct order of operations: PEMDAS = Parentheses, Exponents, Multiplication, Division, Addition, Subtraction 7(1)^3(-2)^-2 Also, what does a negative exponent mean? a^-m = 1/(a^m) Example: p^-5 = 1/p^5

OpenStudy (anonymous):

oh wait... that's not on my answers... answer choices are: 7(-2)^1, 28, 1 3\4 or -7\32

OpenStudy (anonymous):

:( ................

OpenStudy (mathstudent55):

I get 7/4

OpenStudy (mathstudent55):

7(1)^3(-2)^-2 = 7/(-2)^2 = 7/4